Operations Report – March 4th

Crew 223 Operations Report 04/03/2020

SOL: 3

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Tuesday generator:
Off at 5:00am
On at 9:10pm SOC was 84%

Wednesday generator:
Off at 6:00am
On at 9:00pm SOC was XX%

Diesel Reading – 70%

Station Propane Reading – 80%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149694.75 units

Water (static tank) – 334 gallons

Static to Loft Pump used – yes

Water in GreenHab – 180.2 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes
Hours: 149.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover used: no
Hours: 133.6
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: yes
Hours: 139.3
Beginning charge: 100%
Ending charge: 91%
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: Run during the engineer EVA for battery
Oil Added? no
ATV Fuel Used: Nothing to report
Hours the ATVs were Used today: 5 minutes
Notes on ATVs: Nothing to report

HabCar used and why, where? Nothing to report

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 2, 3, 4, 9 were used during the EVA.
Suit 10 is working, and we will use it tomorrow.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: We continue to fix suit 7 but haven’t
finished now. We first found a loose electrical connection at the
connection with the battery at the minus pod. We corrected it. But there
is still a problem with the suit, and we tend for another loose
electrical connection in the control box.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report